# Asymptotic structure of scalar-Maxwell theory at the null boundary

### v1

@@ 4.1 Electric-magnetic duality
$\left\{Q_{\mathrm{el}}[\alpha],Q_{\mathrm{mag}}[\psi]\right\}^{\ast}$ $=$ $C_{\mathrm{EM}}[\alpha,\psi]\,,$ (4.61)
$\left\{Q_{\mathrm{mag}}[\psi_{1}],Q_{\mathrm{mag}}[\psi_{2}]\right\}^{\ast}$ $=$ $0\,.$
@@
The conclusions and discussion of the algebra (4.50) continue being valid also when $\theta\neq 0$.
@@ 5 Electromagnetism coupled to the scalar field
The zero modes are arbitrary on the 2-sphere, and they have determined $r$-dependence as $v_{a}=\left(\frac{v(u,y)}{r},v^{A}(u,y)\right)$. The same $r$dependence will be inherited by the local parameters $\left(\xi,\eta^{A}\right)=\left(\frac{\xi_{(1)}(u,y)}{r},\eta^{A}(u,y)\right)$ of the gauge transformations.
@@
$S[\eta]$ $=$ $\int\mathrm{d}^{3}x\,\eta_{A}\,\Xi^{A}\,.$
While $\alpha$ and $\alpha_{u}$ depend on spacetime coordinates $x^{\mu}$, to ensure that only the first-class constraints will enter in the above generators, we have to impose the following conditions on the local parameters $\eta_{A}$ and $\xi$,
$\xi=\frac{\xi_{(1)}(u,y)}{r}\,,\qquad\eta_{A}=\eta_{A}(u,y)\,.$ (5.33)
As mentioned before, this $r$-dependence has been inherited from the zero modes of the symplectic matrix.
@@
We note that, when $\theta=0$, the new transformations that appear only on the null infinity given by (5.52)–(5.53) are just the shift transformations of the leading components of the fields. The addition of the terms with $\theta\neq 0$ couples non-trivially $\phi$ and the angular components of the electromagnetic field, $A_{A}$.
@@ 6 Discussion
We have examined the asymptotic symmetries of massless theories using the Dirac formalism in the null foliation of Minkowski space-time, providing insights into their structure at null infinity. In general, large gauge charges can also be analyzed in the context of massive theories at timelike infinity, as discussed in QED in [24] and more recently in [25]. It turns out that they also have null infinity contribution [26]. However, the focus of our work is purely on massless theories.
@@
We started by exploring the sector of the theory with $A_{\mu}=0$, corresponding to the scalar field in the Bondi background. We found that the scalar field possesses nontrivial asymptotic symmetries, consisting of an infinite number of Abelian symmetries with the parameters $\xi_{(1)\ell m}$ when expanded in spherical harmonics. Its soft charges coincide with those obtained in [7] via the soft theorem. This intriguing result, that the boundary charges are not related to the gauge symmetries of the bulk, has also been observed in [17]. Nonetheless, the soft scalar charges can be seen as large gauge charges in a dual 2-form field formulation [12].
We then analyzed the sector of the theory with $\phi=\theta=const$, describing Maxwell’s electromagnetic theory supplemented by the topological Pontryagin term. Pure electromagnetism was first discussed using the same formalism in [1], finding asymptotic shift symmetries of the fields $A_{A}(y)$ at the null boundary. They can be interpreted as large gauge transformations of the dual magnetic potential [27], and are related to the magnetic charge $Q_{\mathrm{mag}}$ in the theory [28]. They generate the magnetic soft photon theorem [27] in a similar way as the gauge transformations are related to the electric charge, $Q_{\mathrm{el}}$, and generate the usual Weinberg’s soft photon theorem. Coupled with gravity, non-stationary Maxwell fields contribute to the charges for asymptotic Lorentz symmetries at the null infinity [29].
@@ Acknowledgments
We would like to thank to Miguel Campiglia, Milutin Blagojević, Laurent Freidel, Marc Henneaux, Alfredo Pérez, Marios Petropoulos, Daniele Pranzetti, and Kevin Nguyen for useful discussions. This work has been funded in part by Anillo Grant ANID/ACT210100 Holography and its Applications to High Energy Physics, Quantum Gravity and Condensed Matter Systems and FONDECYT Regular Grants 1221920, 1230492, 1230853, and 1231779.

### v2

@@ 4.1 Electric-magnetic duality
$\left\{Q_{\mathrm{el}}[\alpha_{1}],Q_{\mathrm{el}}[\alpha_{2}]\right\}^{*}$ $=$ $0\,,$
$\left\{Q_{\mathrm{el}}[\alpha],Q_{\mathrm{mag}}[\psi]\right\}^{*}$ $=$ $C_{\mathrm{EM}}[\alpha,\psi]\,,$ (4.61)
$\left\{Q_{\mathrm{mag}}[\psi_{1}],Q_{\mathrm{mag}}[\psi_{2}]\right\}^{*}$ $=$ $0\,.$
@@
It is worth emphasizing that, although the total charge (4.54) remains unchanged for different choices of two independent parameters, this is not necessarily true for the corresponding algebras. It is therefore not clear whether the various Abelian algebras presented in this section would have physically equivalent central extensions.
@@ 5 Electromagnetism coupled to the scalar field
Note that the action (1.1) possesses a global shift symmetry $\phi(x)\to\phi(x)+const$ because the interaction term becomes a topological invariant for a constant scalar.
@@
The zero modes are three arbitrary functions on the 2-sphere. Thus, the kernel of $\Omega_{ss^{\prime}}$ is three-dimensional in the subspace $r=\mathrm{const}.$, where its rank is zero. The $r$-dependence $v_{a}=\left(\frac{v(u,y)}{r},v^{A}(u,y)\right)$ will be inherited by the local parameters $\left(\xi,\eta^{A}\right)=\left(\frac{\xi_{(1)}(u,y)}{r},\eta^{A}(u,y)\right)$, as given by the matrix $\mathbf{Z}$ in (2.10).
@@
$S[\eta]$ $=$ $\int\mathrm{d}^{3}x\,\eta_{A}\,\Xi^{A}\,,$
where, again, the integrals include data at null infinity from $u$ to $u\rightarrow-\infty$. While $\alpha$ and $\alpha_{u}$ depend on spacetime coordinates $x^{\mu}$, to ensure that only the first-class constraints will enter in the above generators, we have to impose the following conditions on the local parameters $\eta_{A}$ and $\xi$,
$\xi=\frac{\xi_{(1)}(y)}{r}\,,\qquad\eta_{A}=\eta_{A}(y)\,.$ (5.33)
As mentioned before, this $r$-dependence has been inherited from the zero modes of the symplectic matrix, while the $u$-independence is the requirement that quotients out trivial asymptotic transformations.
@@
We note that, when $\theta=0$, the new transformations that appear only on the null infinity given by (5.52)–(5.53) are just the shift transformations of the leading components of the fields. Those transformations are asymptotic because they are not symmetries in the bulk of the spacetime. Interestingly, the shift symmetry $\delta\phi=\mathrm{const.}$, which is the true bulk symmetry, has not been detected by this procedure, because the Hamiltonian method sees only symmetries whose parameters depend on some coordinates. The addition of the terms with $\theta\neq 0$ couples non-trivially $\phi$ and the angular components of the electromagnetic field, $A_{A}$.
@@ 6 Discussion
We have examined the asymptotic symmetries of massless theories using the Dirac formalism in the null foliation of Minkowski space-time, providing insights into their structure at null infinity. In general, large gauge charges can also be analyzed in the context of massive theories at timelike infinity, as discussed in QED in [30] and more recently in [31]. It turns out that they also have null infinity contribution [32]. A more general analysis of the role of asymptotically non-trivial gauge transformations, for both massive and massless particles, was discussed in [33]. However, the focus of our work is purely on massless theories.
@@
We started by exploring the sector of the theory with $A_{\mu}=0$, corresponding to the scalar field in the Bondi background. We found that the scalar field possesses nontrivial asymptotic symmetries, consisting of an infinite number of Abelian symmetries with the parameters $\xi_{(1)\ell m}$ when expanded in spherical harmonics. Its soft charges coincide with those obtained in [7] via the soft theorem. This intriguing result, that the boundary charges are not related to the gauge symmetries of the bulk, has also been observed in [29]. Nonetheless, the soft scalar charges can be seen as large gauge charges in a dual 2-form field formulation [12, 34].
We then analyzed the sector of the theory with $\phi=\theta=\mathrm{const.}$, describing Maxwell’s electromagnetic theory supplemented by the topological Pontryagin term. Pure electromagnetism was first discussed using the same formalism in [1], finding asymptotic shift symmetries of the fields $A_{A}(y)$ at the null boundary. They can be interpreted as large gauge transformations of the dual magnetic potential [35], and are related to the magnetic charge $Q_{\mathrm{mag}}$ in the theory [36]. They generate the magnetic soft photon theorem [35] in a similar way as the gauge transformations are related to the electric charge, $Q_{\mathrm{el}}$, and generate the usual Weinberg’s soft photon theorem. In anti-de Sitter space, asymptotic symmetries and their associated charges in electromagnetism have been discussed in [37]. Coupled with gravity, non-stationary Maxwell fields contribute to the charges for asymptotic Lorentz symmetries at the null infinity [38]. In the context of the convolutional double copy, the asymptotic symmetries of Maxwell fields lead to double-copy supertranslations, and the necessity of logarithmic terms in the radial expansion to obtain non-vanishing asymptotic charges [39].
@@
A remaining problem yet to be understood is how the algebra (4.61) depends on the parameter $\theta$. Although the total charge is invariant under $\mathrm{SL}(2,\mathbb{R})$ symmetry, it is not clear whether the corresponding Abelian charge algebra shares this property, that is, whether different choices of parameters lead to equivalent central extensions.
Another, maybe related, question to clarify is the influence of the function $k(y)$ that appears in (4.45) when $\theta=0$ or in (4.52) when $\theta\neq 0$. It can be computed using eq. (4.44) under the integral $\int\mathrm{d}^{2}y\mathrm{d}u\,\partial_{u}$. For example, when $\epsilon=1$, the Hamiltonian $\mathrm{U}(1)$ charge can be written in terms of $A_{A}$ in a gauge-invariant way as
$Q[\alpha]=-\frac{1}{e^{2}}\int\mathrm{d}^{2}y\,\sqrt{\gamma}\,\alpha\nabla^{A}\left[A_{A}(u\rightarrow+\infty,y)-A_{A}(u\rightarrow-\infty,y)\right]\,,$
such that $k(y)=-\frac{1}{e^{2}}\,\sqrt{\gamma}\,\alpha\nabla^{A}A_{(0)A}(u\rightarrow+\infty,y)\,,$ and similarly for $\epsilon=-1$. Due to its field dependence, the quantity $k(y)$ appears as a new canonical variable in the charges. Furthermore, dependence of the electric charge from $\Delta A_{A}=A_{A}(u\rightarrow+\infty,y)-A_{A}(u\rightarrow-\infty,y)$ is a manifestation of the electromagnetic memory effect [42], and it would be interesting to discuss how it would affect central extensions using present formalism.
@@ Acknowledgments
We would like to thank Miguel Campiglia, Milutin Blagojević, Laurent Freidel, Marc Henneaux, Kevin Nguyen, Đorđe Minić, Alfredo Pérez, Marios Petropoulos, and Daniele Pranzetti for useful discussions. We also thank the anonymous referee for their valuable feedback and insightful suggestions, which have helped improve this manuscript. This work has been funded in part by Anillo Grant ANID/ACT210100 Holography and its Applications to High Energy Physics, Quantum Gravity and Condensed Matter Systems and FONDECYT Regular Grants 1221920, 1230492, 1230853, and 1231779.